OpenStudy (anonymous):

1/27 writing it as a power of 2,3 and/or 5.

OpenStudy (anonymous):

i know the answer is 3^-3 but why?

OpenStudy (anonymous):

We know that 3^3=27 correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So \[\frac{1}{27}=\frac{1}{3^3}=3^{−3}\]

OpenStudy (anonymous):

Because \[x^{-n}=\frac{1}{x^n}\]